# Promoting the transition to quantum thinking: development of a secondary school course for addressing knowledge revision, organization, and epistemological challenges

### 31 Jan 2024 v6

@@ III.1 Structuring the learning path
The main source of inspiration and materials for this course has been an educational path for the introduction of QM in the context of polarization developed and evaluated by the PER group of the University of Udine [e.g., 55, 56, 57]. The Udine’s path is focused on the superposition principle and its consequences, starting from the distinction between properties and states. It makes use of hands-on activities with cheap experimental tools (polarizing filters, birefringent crystals), quantitative measurements with light intensity sensors, and of JQM [58], an open-ended environment for computer-simulated experiments on photon polarization. However, the two curricula are different with respect to their design principles, various strategies, physical situations included and sequence. The sequence of activities of our course will be presented in Section IV, and displayed in full in Fig. 15. Here, we describe the the construction of the learning path, showing how the interplay of the first three principles defined its shape.
@@ III.2.2 Superposition
The end of the instructional sequence is represented by the end-of-unit table on superposition. Here we guide students to compare the features of familiar forms of vector superposition (of forces and waves) represented in Fig. 7 with quantum superposition, using prior intuition as a contrast. As a matter of fact, the representational role of this mathematical process in QM is very different from that of the most commonly used forms of superposition in CM. In particular, a consistent interpretation of its referent (only one vector) and its goal (decompose it to obtain info on measurement) is a prerequisite for addressing the quantum notion of interference, which is totally internal to the individual system. From this derives the seemingly contradictory statement of Dirac: “Each photon then interferes only with itself” [64, p. 9]. In such respect, the structure of the frame of Fig. 7 represents a useful template that can be converted into a comparison table, where students are required to identify whether a given statement (e.g.: a goal of superposition is to find the resultant) pertains to one or more forms of superposition (forces, waves, state vectors).
@@ III.4.1 Approaches and research perspectives for running theoretical epistemic practices
While thought experiments may play a variety of roles, we focus on their use as a form of testing procedure. We qualify these procedures as humble thought experiments, because they are not meant to achieve the purposes of historically significant thought experiments (e.g., Einstein’s elevator); yet, their structure corresponds to that described by Gilbert and Reiner [50] for a thought experiment, and their conduction may be within the reach of secondary school students. In order to design and analyze tasks in which students are actively engaged in running a thought experiment, we refer to the ISLE learning framework [51], in which the phases of testing experiments are clearly identified and associated with scientific abilities. These abilities are described in rubrics introduced in Etkina et al. [68], which are to be used as self-assessment instruments. For scientific abilities related to testing experiments, see Etkina [51, Appendix B]. While the rubrics have been developed as a support in the design and running of experiments in the lab, they may be easily adapted to activities in which the instructor
@@ III.4.4 Thought experiment: Superposition as statistical mixture of component states?
The scientific abilities that are associated with the conduction of the thought experiment are again a subset of those described by Etkina, and are expressed as in her article [51, Appendix B]:
@@ III.5.1 Impact of the interpretive choices on the coherence of the design
While we feel that this perspective can be perceived as plausible by students, who can make a connection with already familiar classical fields (especially in the case of a photon), ascribing a quantum field ontology to physical systems is a controversial operation [e.g., 70, 71, pp. 133-135]. For this reason, we adhere to a cautious approach, suggesting students to model the system as a “field of actual and potential properties.” This expression means that the field describes the system in terms of properties it possesses (e.g., one might be a property of a spin component) and of “potential properties that can possibly be actualized […] through measurement processes” [72]. For more information on the concept of potential property and its transition to actuality, see also C. J. Isham [73]. Based on the examination of the “which-way" experiment, students are led to identify two further elements of revision in the concept of field: differently from a classical field, a quantum one displays a punctual interaction with detectors (we can identify the detector with which the interaction takes place), but this interaction affects the entire field at the same instant, i.e., in a non-local way.
@@ III.5.2 Structuring the discussion of epistemological themes
Format, content and placement of the activities on the foundational debate need not only be instrumental to the implementation of the Epistemological Principle, but also compatible with the educational level of the student population at hand, the structure of the learning path, and the duration of the course (12 hours). The chosen format consists in a short introductory lecture given by the instructor, followed by a whole class discussion of the topic, which can be supported by pre-class reading assignments. This structure is in accordance with the method used to conduct worksheet-based activities (see Section IV.3), which represent the common thread of the course, and is adapted to the complexity of foundational topics by extending the time for the whole class debate, that needs to be at center stage, and by providing multiple forms of support. The texts are preferably selected among those works of leading scientists whose understanding does not require sophisticated mathematical or physical knowledge [e.g., excerpts from 74, 75]. Since the first three units of the course concern preparation, measurement, and their formalism, while propagation, wave-particle duality, and entanglement are addressed in Unit 4, it is natural to discuss first the debates on indefiniteness, uncertainty, and completeness.
@@
After the conceptual and mathematical examination of the polarization entanglement of two photons produced by parametric down-conversion, students are presented with the problem of non-locality, which is discussed only at a qualitative level. This discussion allows us to emphasize the importance of the foundational debate in the development of scientific knowledge. In the words of Alain Aspect, “there was a lesson to be drawn: questioning the ‘orthodox’ views, including the famous ‘Copenhagen interpretation’, might lead to an improved understanding of the quantum mechanics formalism, even though that formalism remained impeccably accurate” [75, p. xix]. As regards technological development, it is possible to illustrate to students that a deeper understanding of entanglement is at the root of a second quantum revolution that is now unfolding [e.g., 77], and that John Bell has been its prophet [75].
@@ V.1 Design-Based Research: data collection and analysis
Since the sample changed from experiments to experiment, in order to improve readability and to enable comparison, the rates of answers as regards both correctness and student reasoning are reported by means of percentages.
@@ V.2.1 Measurement
Path of refinement. In the initial version of the course (2014 Summer School), the introduction of measurement was scheduled right after an extensive work with polarizing filters, both at a macroscopic level and in terms of photons (see Section V.4.2). Therefore, we used, as a measurement device, a filter followed by a counter. However, the design experiment showed that only a small number of students spontaneously interpreted the absorption of a photon by a filter in terms of acquisition or retention of a property in the direction perpendicular to its axis and, as a consequence, that most did not recognize the conditional nature of measurement. The addition of an iconic support in 2015 (two diagrams depicting the possible transitions of the initial property in case of transmission and of absorption) was not effective. In the stochastic case, only a minority interpreted absorption in terms of acquisition of a property in the direction perpendicular to the axis. Some students even wondered how the whole situation could be described as a measurement, and not as “just a weird interaction altering the property”. Given the need to provide students with a context where the results of the interaction with the measurement device can always be visualized as the acquisition/retention of outcome-properties, we resolved to replace the filter with a birefringent crystal and two counters.
@@
Given the high rate of success (see the progression in Fig. 19), in the following design experiments the items were converted into oral questions, since we intended to focus on the refinement of later parts of the course.
@@ V.2.2 Superposition
Path of refinement. The first version of the table was administered in the 2017 Summer School. At that time, the course focused only on polarization, since the context of the hydrogen-like atom was still in development. The table included four statements and asked students to assess whether each of them applied to the superposition of forces, waves, and state vectors, and to explain their reasoning in the last context. The statements concerned the aspects of superposition examined in Section III.2.2: referent, goal, procedure, and constraints. The task on the referent did not address the number of entities involved, but was aimed to reinforce the discrimination between the abstract nature of state vectors and the physical nature of the classical ones, which proved to be a deep issue in the introduction of the state vector (see Section III.4.5). While a majority of students consistently answered this item and the item on the constraints, the tasks on the goal and the procedure gave very poor results, due to difficulties with the number of entities involved in quantum superposition. This outcome prompted us to revise both the previous activities of Unit 3 and the table. In order to support students in recognizing that quantum superposition refers only to one entity - the state of the system -, we added an activity relying on embodied cognition (see Zuccarini and Malgieri [31]). In that year, we also added a more general context, i.e., the superposition of eigenstates of the hydrogen-like atom in terms of quantum numbers.
@@
In the first row, we included a statement on the number of entities involved, which was considered as preparatory to the discussion of the goal, placed in the third row. We left unchanged the two productive items on the nature of the vector and on the constraints. The item on the procedure was removed, since both the first and the third item already invited a discussion of this aspect.
@@
Figure 21: Boxes on quantum superposition: correct answers with consistent explanations in 2017 and 2018
A strong improvement is evident in the answers to the first and the third item, with 87% of the students providing consistent explanations. As regards the number of entities involved, 53% did not limit themselves to state that quantum superposition concerns one entity, but interpreted the process as a decomposition of this entity. As to the goal, beside recognizing that determining the resultant is not among the goals of this form of superposition, most students identified its objectives as “calculating [transition] probabilities” (37%) or “decomposing the state vector” (30%). In both design experiments, more than 65% consistently answered the other two items on the nature of the vector and on the constraints, with a slight decrease in performance in 2018. Students justified their answer about the nature of the vector by adding consistent explanations such as “state vectors are unit vectors with no measurement units”, “they express probability”. As regards the statement on the constraints, we need to take into account the greater complexity introduced by the superposition of states of the hydrogen-like atom (where, for bound states, we can have a countably infinite number of components, all orthogonal to one another). Most students discussed the statement by referring to the context of polarization, e.g. “In polarization, for instance, we can have up to two components $\rightarrow$ two values” (47%). However, some students gave global explanations by connecting the number of vectors to the spectrum of the measured observable: “No, it depends on the number of values that a property can assume” (13%), which is true in the absence of degeneracy.
@@ V.3.1 Introducing the relations between properties
Figure 23: Results of the task. In this graph and in the next, P.S. means percentage of students.
@@
Figure 24: Comparison of the results of the task number $1$: identity; 2: mutual unacquirability; 3: incompatibility; 4: compatibility.
@@ V.3.2 Extending the use of the relations between properties to other contexts
While in both experiments, more than half of the students consistently answered all of the questions, their results were generally worse than that obtained in the first block, with summer school students outperforming regular classrooms. As to the latter, we still get a very high percentage of consistent answers to F1.1 (85%) and F2.1 (81%), while 60% consistently answered F2.2. Question F2.1 turned out to be the most difficult for them, with 53% of consistent answers. Many students did not connect the situation described in the item to the relation of compatibility Others said that if the properties are compatible, nothing changes in measurement In question F2.2, instead, student explanations were quite similar to those reported for the fourth item on $x$ and $v$. In general, the physical situations presented in question F1.2 and F2.2 were totally new to students, who had never encountered phenomena related to compatible observables in previous parts of the course. Here, the carefully selected and motivated students of the summer school have been quicker to respond to the challenge posed by the new context.
@@ VI Conclusions
Learning quantum mechanics involves overcoming manyfold challenges, among which knowledge revision, fragmentation, and the construction of a personally plausible picture of the quantum model. Extensive investigations of cognitive and epistemological challenges, as well of the use of student resources and intuition in the discussion of introductory science topics have been conducted in the framework of conceptual change [5, 6], leading to the development of effective approaches to science teaching [28]. Also in the case of QM, various researchers have considered the problem of teaching the subject as the design of strategies to effectively promote a conceptual change in individual learners [e.g., 22, 27, 26, 9]. In general, the progress from a classical to a quantum perspective requires changes in knowledge structures about a scientific theory, which have been developed as a result of instruction. Since educational models of conceptual change have been introduced to account for a different kind of change (the transition from naïve to scientific knowledge), there is a need to revise the framework, examining the differences involved in learning a successive theory. In addition, while studies on conceptual change are undergoing a systemic turn [28], there is a lack of educational proposals on QM that coordinate multiple interacting aspects at different levels of analysis. Last, in accordance with the shift in focus from difficulties to resources [18], there is a need to identify the links between student intuition and strategies for learning QM.
In this article, we describe the development of a course for secondary school that is based on an examination of conceptual change in the transition from CM to QM, and integrates cognitive and epistemological aspects including the adoption of epistemic practices of theoretical physicists, and a structured approach to interpretive themes. The aim is to help fill the gaps in analysis and design, addressing all the challenges mentioned at the beginning of this section, and taking advantage of available knowledge elements and intuition in the task.

### 20 Feb 2024 v8

@@ III.1 Structuring the learning path
The main source of inspiration and materials for this course has been an educational path for the introduction of QM in the context of polarization developed and evaluated by the physics education research group of the University of Udine [e.g., 54, 55, 56]. The Udine’s path is focused on the superposition principle and its consequences, starting from the distinction between properties and states. It makes use of hands-on activities with cheap experimental tools (polarizing filters, birefringent crystals), quantitative measurements with light intensity sensors, and of JQM [57], an open-ended environment for computer-simulated experiments on photon polarization. However, the two curricula are different with respect to their design principles, various strategies, physical situations included and sequence. The sequence of activities of our course will be presented in Section IV, and displayed in full in Fig. 15. Here, we describe the the construction of the learning path, showing how the interplay of the first three principles defined its shape.
@@ III.2.2 Superposition
The end of the instructional sequence is represented by the end-of-unit table on superposition. Here we guide students to compare the features of familiar forms of vector superposition (of forces and waves) represented in Fig. 7 with quantum superposition, using prior intuition as a contrast. As a matter of fact, the representational role of this mathematical process in QM is very different from that of the most commonly used forms of superposition in CM. In particular, a consistent interpretation of its referent (only one vector) and its goal (decompose it to obtain info on measurement) is a prerequisite for addressing the quantum notion of interference, which is totally internal to the individual system. From this derives the seemingly contradictory statement of Dirac: “Each photon then interferes only with itself” [63, p. 9]. In such respect, the structure of the frame of Fig. 7 represents a useful template that can be converted into a comparison table, where students are required to identify whether a given statement (e.g.: a goal of superposition is to find the resultant) pertains to one or more forms of superposition (forces, waves, state vectors).
@@ III.4.1 Approaches and research perspectives for running theoretical epistemic practices
While thought experiments may play a variety of roles, we focus on their use as a form of testing procedure. We qualify these procedures as humble thought experiments, because they are not meant to achieve the purposes of historically significant thought experiments (e.g., Einstein’s elevator); yet, their structure corresponds to that described by Gilbert and Reiner [49] for a thought experiment, and their conduction may be within the reach of secondary school students. In order to design and analyze tasks in which students are actively engaged in running a thought experiment, we refer to the ISLE learning framework [50], in which the phases of testing experiments are clearly identified and associated with scientific abilities. These abilities are described in rubrics introduced in Etkina et al. [67], which are to be used as self-assessment instruments. For scientific abilities related to testing experiments, see Etkina [50, Appendix B]. While the rubrics have been developed as a support in the design and running of experiments in the lab, they may be easily adapted to activities in which the instructor
@@ III.4.4 Thought experiment: Superposition as statistical mixture of component states?
The scientific abilities that are associated with the conduction of the thought experiment are again a subset of those described by Etkina, and are expressed as in her article [50, Appendix B]:
@@ III.5.1 Impact of the interpretive choices on the coherence of the design
While we feel that this perspective can be perceived as plausible by students, who can make a connection with already familiar classical fields (especially in the case of a photon), ascribing a quantum field ontology to physical systems is a controversial operation [e.g., 69, 70, pp. 133-135]. For this reason, we adhere to a cautious approach, suggesting students to model the system as a “field of actual and potential properties.” This expression means that the field describes the system in terms of properties it possesses (e.g., one might be a property of a spin component) and of “potential properties that can possibly be actualized […] through measurement processes” [71]. For more information on the concept of potential property and its transition to actuality, see also C. J. Isham [72]. Based on the examination of the “which-way" experiment, students are led to identify two further elements of revision in the concept of field: differently from a classical field, a quantum one displays a punctual interaction with detectors (we can identify the detector with which the interaction takes place), but this interaction affects the entire field at the same instant, i.e., in a non-local way.
@@ III.5.2 Structuring the discussion of epistemological themes
Format, content and placement of the activities on the foundational debate need not only be instrumental to the implementation of the Epistemological Principle, but also compatible with the educational level of the student population at hand, the structure of the learning path, and the duration of the course (12 hours). The chosen format consists in a short introductory lecture given by the instructor, followed by a whole class discussion of the topic, which can be supported by pre-class reading assignments. This structure is in accordance with the method used to conduct worksheet-based activities (see Section IV.3), which represent the common thread of the course, and is adapted to the complexity of foundational topics by extending the time for the whole class debate, that needs to be at center stage, and by providing multiple forms of support. The texts are preferably selected among those works of leading scientists whose understanding does not require sophisticated mathematical or physical knowledge [e.g., excerpts from 73, 74]. Since the first three units of the course concern preparation, measurement, and their formalism, while propagation, wave-particle duality, and entanglement are addressed in Unit 4, it is natural to discuss first the debates on indefiniteness, uncertainty, and completeness.
@@
After the conceptual and mathematical examination of the polarization entanglement of two photons produced by parametric down-conversion, students are presented with the problem of non-locality, which is discussed only at a qualitative level. This discussion allows us to emphasize the importance of the foundational debate in the development of scientific knowledge. In the words of Alain Aspect, “there was a lesson to be drawn: questioning the ‘orthodox’ views, including the famous ‘Copenhagen interpretation’, might lead to an improved understanding of the quantum mechanics formalism, even though that formalism remained impeccably accurate” [74, p. xix]. As regards technological development, it is possible to illustrate to students that a deeper understanding of entanglement is at the root of a second quantum revolution that is now unfolding [e.g., 76], and that John Bell has been its prophet [74].
@@ V.1 Design-Based Research: data collection and analysis
Since the sample changed from experiments to experiment, in order to improve readability and to enable comparison, the rates of answers as regards both correctness and student reasoning are reported in percentages.
@@ V.2.1 Measurement
Path of refinement. In the initial version of the course (2014 Summer School), the introduction of measurement was scheduled right after an extensive work with polarizing filters, both at a macroscopic level and in terms of photons (see Section V.4.2). Therefore, we used, as a measurement device, a filter followed by a counter. However, the design experiment showed that only a small number of students spontaneously interpreted the absorption of a photon by a filter in terms of acquisition or retention of a property in the direction perpendicular to its axis and, as a consequence, that most did not recognize the conditional nature of measurement. The addition of an iconic support in 2015 (two diagrams depicting the possible transitions of the initial property in case of transmission and of absorption) was not effective. In the stochastic case, only a minority interpreted absorption as a result of the acquisition of a property. Some students even wondered how the situation could be described as a measurement, and not as “just a weird interaction altering the property”. Given the need to provide students with a context where the results of the interaction with the measurement device can always be visualized as the acquisition/retention of outcome-properties, we resolved to replace the filter with a birefringent crystal and two counters.
@@
Given the high rate of success (see the progression in Fig. 19), in the following design experiments the items were converted into oral questions.
@@ V.2.2 Superposition
Path of refinement. The first version of the table was administered in the 2017 Summer School. At that time, the course focused only on polarization. The table included four statements and asked students to assess whether each of them applied to the superposition of forces, waves, and state vectors, and to explain their reasoning in the last context. The statements concerned the aspects of superposition examined in Section III.2.2: referent, goal, procedure, and constraints. The task on the referent did not address the number of entities involved, but was aimed to reinforce the discrimination between the abstract nature of state vectors and the physical nature of the classical ones, which proved to be a deep issue in the introduction of the state vector (see Section III.4.5). While a majority of students consistently answered this item and the item on the constraints, the tasks on the goal and the procedure gave very poor results, due to difficulties with the number of entities involved in quantum superposition. This outcome prompted us to revise both the previous activities of Unit 3 and the table. In order to support students in recognizing that quantum superposition refers only to one entity - the state of the system -, we added an activity relying on embodied cognition (see Zuccarini and Malgieri [30]). We also added a more general context: the superposition of eigenstates of the hydrogen-like atom in terms of quantum numbers.
@@
In the first row, we included a statement on the number of entities involved, which was considered as preparatory to the discussion of the goal. We left unchanged the two productive items on the nature of the vector and on the constraints. The item on the procedure was removed, since both the first and the third item already invited a discussion of this aspect.
@@
Figure 21: Answers with correct explanation in 2017 and 2018.
A strong improvement is evident in the answers to the first and the third item, with 87% of the students providing consistent explanations. As regards the number of entities involved, 53% did not limit themselves to state that quantum superposition concerns one entity, but interpreted the process as a decomposition of this entity. As to the goal, beside recognizing that determining the resultant is not among the goals of this form of superposition, most students identified its objectives as “calculating [transition] probabilities” (37%) or “decomposing the state vector” (30%). In both design experiments, more than 65% consistently answered the other two items on the nature of the vector and on the constraints, with a slight decrease in performance in 2018. Students justified their answer about the nature of the vector by adding consistent explanations such as “state vectors are unit vectors with no measurement units”, “they express probability”. In the statement on the constraints, we need to consider the greater complexity introduced by the context of the hydrogen-like atom (where, for bound states, we can have a countably infinite number of components, all orthogonal to one another). Most students discussed the statement only in the context of polarization, e.g. “In polarization, for instance, we can have up to two components $\rightarrow$ two values” (47%). However, some students gave global explanations by connecting the number of vectors to the spectrum of the measured observable: “No, it depends on the number of values that a property can assume” (13%), which is true in the absence of degeneracy.
@@ V.3.1 Introducing the relations between properties
Figure 23: Identifying the relations. Results of the task. In this graph and in the next, P.S. means percentage of students.
@@
Figure 24: Identifying the relations. Comparison of consistent + partial answers. 1: Identity; 2: Unacquirability; 3: Incompatibility; 4: Compatibility.
@@ V.3.2 Extending the use of the relations between properties to other contexts
While in both experiments, more than half of the students consistently answered all of the questions, their results were generally worse than that obtained in the first block, with summer school students outperforming regular classrooms. As to the latter, we still get a very high percentage of consistent answers to F1.1 (85%) and F2.1 (81%), while 60% consistently answered F2.2. Question F1.2 turned out to be the most difficult for them, with 53% of consistent answers. Many students did not connect the situation described in the item with the relation of compatibility Others said that if the properties are compatible, nothing changes in measurement In question F2.2, an analogous but more complex version of the fourth item on $x$ and $v$, student explanations were quite similar to those reported for the latter. In general, the physical situations presented in question F1.2 and F2.2 were totally new to students, who had never encountered phenomena related to compatible observables in previous parts of the course. Here, the carefully selected and motivated students of the summer school were quicker to respond to the challenge posed by the new context.
@@ VI Conclusions
Learning quantum mechanics involves overcoming manyfold challenges, among which knowledge revision, fragmentation, and the construction of a personally plausible picture of the quantum model. Extensive investigations of cognitive and epistemological challenges, as well of the use of student resources and intuition in the discussion of introductory science topics have been conducted in the framework of CC [5, 6], leading to the development of effective approaches to science teaching [27]. Also in the case of QM, various researchers have considered the problem of teaching the subject as the design of strategies to effectively promote a CC in individual learners [e.g., 21, 26, 25, 9]. In general, the progress from a classical to a quantum perspective requires changes in knowledge structures about a scientific theory, which have been developed as a result of instruction. Since educational models of CC have been introduced to account for a different kind of change (the transition from naïve to scientific knowledge), there is a need to revise the framework, examining the differences involved in learning a successive theory. In addition, while studies on CC are undergoing a systemic turn [27], there is a lack of educational proposals on QM that coordinate multiple interacting aspects at different levels of analysis. Last, in accordance with the shift in focus from difficulties to resources [17], there is a need to identify the links between student intuition and strategies for learning QM.
In this article, we describe the development of a course for secondary school that is based on an examination of CC in the transition from CM to QM, and integrates cognitive and epistemological aspects including the adoption of epistemic practices of theoretical physicists, and a structured approach to interpretive themes. The aim is to help fill the gaps in analysis and design, addressing all the challenges mentioned at the beginning of this section, and taking advantage of available knowledge elements and intuition in the task.